OpenStudy (anonymous):

need help : Factor completely x^2+7xy-144y^2

OpenStudy (anonymous):

What are the factors of 144 1 * 144 2 * 72 3 * 46 4 * 36 6 * 24 8 * 18 9 * 16 12 * 12 Which pair subtracts to get you 7?

OpenStudy (nikvist):

\[x^2+7xy-144y^2=x^2+16xy-9xy-144y^2=\] \[=x(x+16y)-9y(x+16y)=(x+16y)(x-9y)\]